# Domain-adaptive Post-training of Large Language Models: An Empirical Study with Llama for Finance [Yifei: A more attractive/informative title?]

### v2

@@ Abstract
Zixuan Ke, Yifei Ming, Xuan-Phi Nguyen, Caiming Xiong and Shafiq Joty Salesforce AI Research {zixuan.ke,yifei.ming,xnguyen,cxiong,sjoty}@salesforce.com [Uncaptioned image] Project Page: https://github.com/SalesforceAIResearch/FinDAP [Uncaptioned image] Datasets: https://huggingface.co/datasets/Salesforce/FinEval
@@ 1 Introduction
In this work, we introduce FinDaP (Figure 1), a novel finance-specific framework designed to incorporated all these factors in domain-specific post-training. To our knowledge, none of the prior studies consider all of them to provide a principled guidance on domain-adaptive post-training. FinDaP integrates four key components: (1) FinCap, a set of core capabilities required for the domain expert LLM, including domain concepts, tasks, instruction following and reasoning; (2) FinRec, a training recipe that jointly performs CPT and IT, and subsequently conducts PA, balancing trade-offs across these stages to mitigate forgetting and improve task generalization. It also proposes to use mixture of in-domain and general domain data in the data recipe, alongside a novel preference alignment method for improving reasoning capability that constructs data using the preference signal in reasoning steps, Stepwise Corrective Preference (SCP) and final answer, Final Answer Preference (FAP); (3) FinTrain, a curated set of training datasets implementing FinRec, which carefully balances quality and diversity; and (4) FinEval, a comprehensive evaluation framework covering a wide range of tasks, including reasoning tasks assessed through CoT.
We apply FinDaP on the instruction-tuned Llama3-8b-instruct (LLaMA, 2024). Our best performing recipe yields Llama-Fin that outperforms all considered baselines, including large open models at the 70B scale and proprietary models like GPT-4o, on tasks that are similar (yet unseen) to the training data. Even on novel tasks that were never encountered in training, Llama-Fin remains competitive and consistently outperforms its base model across all identified capabilities.
@@ 2 Related Work
Finance LLM Capabilities Recipe Evaluation
@@
AdaptLLM Concept CPT CPT: Financial text + heuristic QAs constructed from the text Financial + Classification tasks + Direct answer
PIXIU Task IT IT: Financial tasks Financial + Classification tasks + Direct answer
FinLLM Concept, Task CPT → IT CPT: Financial text + Fineweb; IT: Filtered Financial tasks Financial + Classification tasks + Direct answer
FinTral Concept, Task CPT → IT → PA CPT: Financial text; IT: Financial tasks; PA: Outcome signal only Financial + Classification tasks + Direct answer
@@
Finance LLMs Table 1 summarizes popular finance-specific LLMs developed through domain-adaptive post-training. AdaptLLM (Cheng et al., 2024) focuses on CPT and constructs heuristic QA tasks from raw text, but it considers only five financial end tasks. PIXIU (Xie et al., 2023) focus on instruction-followingby creating a financial instruction-tuning dataset from diverse open financial tasks and designing a benchmark with nine end tasks for evaluation. FinLLM (Xie et al., 2024a) extends post-training across multiple stages, first performing CPT, then IT, and incorporating multi-modal capabilities via IT. It includes some general-domain data (e.g., fineWeb) but does not explore its impact systematically. Following this line, FinTral (Bhatia et al., 2024) follows this approach but is the only open FinLLM to include PA, where preference labels were given by GPT-4 on the final outcome, considering only coarse-grained signals. It also introduces multi-modality via IT and integrates tool use and retrieval in PA training. Additionally, Palmyra-Fin (Writer, 2024), a recent state-of-the-art FinLLM, reports high performance on finance tasks, particularly CFA exams [^1] https://www.cfainstitute.org/programs/cfa-program/exam, but its training recipe remains undisclosed.
Comparing to FinDaP, none of these models explicitly identify desirable capabilities as we do with FinCap, nor do they systematically explore trade-offs between CPT, IT and PA to develop a more effective training recipe. They also do not incorporate fine-grained process signals in PA to improve reasoning, as we do in FinRec. Additionally, their evaluations lack the broader range of tasks, methods, and similarities, including reasoning tasks and chain-of-thought evaluations, that we adopt in FinEval. Finally, unlike Palmyra-Fin, Llama-Fin is fully open-source, ensuring complete transparency in its training recipe, datasets, and evaluation methods, while achieving SoTA in its size category.
@@ 3.1 Core Capabilities (FinCap)
We began by conducting a comprehensive survey of existing work and consulting two financial domain experts—a banking industry advisor and a financial industry product manager. Through this process, we identified four key fundamental capabilities essential for a finance-focused LLM: understanding domain-specific concepts to process financial language accurately, performing domain-specific tasks to solve real-world problems, reasoning effectively to analyze complex financial data, and following instructions to interact naturally in practical applications. These capabilities are deeply interconnected—reasoning depends on conceptual knowledge, while instruction-following ensures effective communication.
@@ 3.2.1 Model Recipe
Previous studies often de facto treat domain-adaptive post-training as a sequential process involving, or partially involving, CPT, IT, and PA. However, our experiments with LLaMA3-8B-Inst show key trade-offs among these stages (Appendix §B). While CPT is effective at introducing domain concepts, it often leads to significant forgetting of general concepts and instruction-following capabilities. In contrast, IT strengthens instruction-following capabilities and introduces domain-specific tasks with minimal forgetting. IT alone however struggles with task generalization. PA is effective for learning reasoning but depends heavily on high-quality preference data, which can be difficult to synthesize. To address these limitations, we propose a joint CPT+IT approach, resulting in CPT+IT checkpoint. Subsequently, PA is performed with a novel trajectory collection method that provides fine-grained supervision signals.
Joint continual pre-training and instruction-tuning (CPT + IT)
@@
In this stage, the goal is to learn domain-specific knowledge while maintaining general capabilities, such as instruction-following. It is well known that CPT can adapt the LLM to learn domain-specific concepts while IT can help learn the domain-specific and instruction-following tasks Ke et al. (2022); Wei et al. (2022). Typically CPT involves next-token prediction without masking any context tokens, and IT involves next-token prediction with instructions masked out; thus training them sequentially from an instruction-tuned LLM naturally leads to forgetting general capabilities, including instruction-following. Intuitively, if the loss function incorporates both CPT and IT, forgetting can be largely mitigated Scialom et al. (2022). [^2] This is akin to ‘replay’ method in continual learning. To achieve this, we mix CPT and IT data, effectively performing joint optimization, as the only difference between the two is whether the instruction is masked. This approach also facilitates knowledge transfer, as CPT helps the model learn domain knowledge, which can be leveraged by IT training. More importantly, since concepts learned from CPT are often inherently more generalizable due to the shared nature of concepts across tasks, jointly training CPT and IT can improve generalization without require exposure to a diverse range of tasks Wei et al. (2022), which is often impractical in certain domains, particularly long-tail ones. Since CPT datasets are typically much larger than IT datasets, we downsample CPT data to match the size of IT data, allowing for effective joint training.
Improving reasoning with preference alignment
@@
CPT+IT improves capabilities such as in general and domain-specific concepts, tasks and IF/Chat. However, we find that the resulting model lacks in its reasoning capability, especially when it comes to complex reasoning like solving problems in CFA exams, where it is important to make each reasoning step correct. We use PA for this, which trains the model to assign higher probability mass to better generations, and has been shown to be effective in enhancing LLM reasoning capabilities (Lambert et al., 2024; Jiao et al., 2024). Specifically, we employ Direct Preference Optimization or DPO (Rafailov et al., 2023), which allows the model to learn from both positive and negative examples, providing a richer learning signal compared to SFT. We synthetically generate such data from the on-policy model, i.e., the jointly trained CPT+IT checkpoint, as it has shown the strongest performance in previous experiments. We propose a novel trajectory collection method that provides fine-grained step-level supervision signals (§3.2.2)
@@ 3.2.2 Data Recipe
Most existing finance LLMs rely exclusively on in-domain data in post-training with the exception of FinLLM, which uses general domain data in CPT (see Table 1). Intuitively, this exclusive reliance on in-domain data can lead to forgetting of general knowledge in the original pre-trained LLM. To understand how the forgetting happens across different stages, we conduct ablation by constructing three versions of data for each training stage: in-domain, general-domain, and a mixture of both. Experiments (Appendix §B) show that the impact of forgetting decreases progressively from CPT to IT to PA, with CPT experiencing the most severe forgetting and PA the least. Guided by these findings, we adopt the mixture of in- and general-domain data for CPT+IT training to maximize both specialization and retention of essential general knowledge.
Preference data construction for reasoning
@@
Figure 2: An overview of the proposed final answer preference (FAP) and stepwise corrective preference (SCP). In FAP, we collect trajectories from the GenRM by evaluating the entire solution (thus question marks in the intermediate steps). In SCP, we collect trajectories from the GenRM, by identifying and correcting the first erroneous step.
@@
Existing training methods to improve reasoning primarily rely on outcome-based rewards, which provide sparse supervision and do not guide intermediate reasoning steps. At the same time, stepwise reward models can be computationally expensive if applied at every step. To strike a balance, we employ a Generative Reward Model (GenRM) and design Final Answer Preference (FAP) to efficiently collect preference signals at the final answer level (outcome reward), while also collecting Stepwise Corrective Preference (SCP) at the reasoning step level (process reward) by asking the GenRM to identify and correct the first erroneous step. By combining these two complementary strategies, our PA provides stronger supervision signals for reasoning improvements while maintaining efficiency, making it particularly suitable for domains like finance, where both accuracy and efficiency are critical. Figure 2 illustrates the proposed method.
@@
$\bullet$ Final Answer Preference (FAP). Given a prompt and a model generated solution, we prompt the GenRM to give a holistic judgment for the entire solution using a single ‘Yes’ or ‘No’ token. We then use the correct solutions as chosen samples and the incorrect solutions as rejected samples.
@@
$\bullet$ Stepwise Corrective Preference (SCP). Since reasoning could be complex (e.g., CFA exams) and process rewards have been shown to be more effective in such cases (Lightman et al., 2024), we further leverage the GenRM to provide step-level signals. Instead of requesting rewards at each step, which has been shown to be unnecessary in Lightman et al. (2024); Luo et al. (2024), we prompt the GenRM to identify the first erroneous step and ask it to provide a correction for that step. Using this correction, we construct a preference data sample. The input prompt is formed by concatenating the original question, the candidate reasoning steps up to the first error, and a follow-up question framed as “What is the next step?”. The chosen response of this preference sample is the newly-obtained corrected step, while the original first erroneous step is deemed as rejected response. This approach produces trajectories that focus on predicting the correct next step given a reasoning prefix, rather than requiring a prediction of the entire reasoning trajectory, as in FAP. Additional details about this approach are provided in Appendix G.
@@ 3.3 FinDaP Training Data (FinTrain)
In FinTrain, we carefully balance the trade-off between quality and quantity of the training data at each stage. Specifically, in CPT, we leverage available general-domain supervised data, such as NaturalInstrution Mishra et al. (2022) and Aqua Ling et al. (2017). Since such data has been carefully curated and cleaned for labeling, they maintain good quality. Furthermore, because of the presence of the supervised responses, they can be helpful in mitigating the forgetting issue (Scialom et al., 2022). For quantity and diversity, which is essential for learning new domain knowledge during CPT, we collect large-scale, diverse data from relevant sources, including 70 financial websites and books covering 12 financial topics, e.g., CFA exam preparation materials. We further use a strong LLM to filter out low-quality tasks based on an additive scale prompt Yuan et al. (2024). This results in a CPT dataset of approximately 8B tokens.
For IT, to promote diversity, we conduct a broad survey and source general, financial, instruction-following, and reasoning datasets from public datasets. We also include large open QA datasets like FinQA Chen et al. (2021). To ensure quality, we prioritize datasets that shown to perform well in the literature, e.g., UltraChat Ding et al. (2023). We also incorporate exercises or demonstrations from books that often contain human-written CoT. The final IT dataset consists of approximately 3M prompts.
For PA, we use CFA preparation materials as a representative source for in-domain reasoning as they cover diverse financial scenarios, em-phasize complex reasoning, and, most importantly, are derived from real-world exams. We construct preference data with FAP and SCP introduced in §3.2.2. The final PA dataset consists of approximately 32K prompts. Additional details are given in Appendix D.
@@ 3.4 FinDaP Evaluation (FinEval)
Our evaluation framework FinEval is designed to systematically assess model performance across unseen tasks. Unlike prior studies that rely on a narrow set of domain-specific tasks, FinEval categorizes tasks by similarity (similar vs. novel tasks), domain specificity (general vs. domain-specific vs. reasoning tasks), and evaluation methods (direct-answer vs. chain-of-thought). By structuring evaluations along these dimensions, In total, FinEval consists of 35 tasks and can serve as a comprehensive benchmark for the expected capabilities going beyond simple task-based evaluation. We provide details about the evaluation tasks and methods in Appendix E.

### v4

# Demystifying Domain-adaptive Post-training for Financial LLMs

@@ Abstract
Zixuan Ke, Yifei Ming, Xuan-Phi Nguyen, Caiming Xiong, Shafiq Joty Salesforce AI Research {zixuan.ke,cxiong,sjoty}@salesforce.com [Uncaptioned image] Project Page: https://vincent950129.github.io/adapt-llm/ [Uncaptioned image] Llama-Fin-8b: https://huggingface.co/Salesforce/Llama-Fin-8b [Uncaptioned image] FinTrain: https://huggingface.co/datasets/Salesforce/FinTrain [Uncaptioned image] FinEval: https://huggingface.co/datasets/Salesforce/FinEval [Uncaptioned image] Code: https://github.com/SalesforceAIResearch/FinDAP
@@ 1 Introduction
In this work, we introduce FinDaP (Figure 1), a novel finance-specific framework designed to incorporate all these factors in domain-specific post-training. To our knowledge, none of the prior studies consider all of them to provide a principled guidance on domain-adaptive post-training. FinDaP integrates four key components: (1) FinCap, a set of core capabilities required for the domain expert LLM, derived from a systematic review of prior literature and input from domain experts in finance. These include domain concepts, tasks, instruction following and reasoning; (2) FinRec, a training recipe that jointly performs CPT and IT, and subsequently conducts PA, balancing trade-offs across these stages to mitigate forgetting and improve task generalization. It also proposes to use mixture of in-domain and general domain data in the data recipe, alongside a novel preference alignment method for improving reasoning capability that constructs data using the preference signal in reasoning steps, Stepwise Corrective Preference (SCP), and final answer, Final Answer Preference (FAP); (3) FinTrain, a curated set of training datasets implementing FinRec, which carefully balances quality and diversity; and (4) FinEval, a comprehensive evaluation framework covering a wide range of tasks, including reasoning tasks assessed through CoT. [^1] We will open-source the data, checkpoint, code, leaderboard for all components upon acceptance.
We apply FinDaP on the instruction-tuned Llama3-8b-instruct (LLaMA, 2024). Our best performing recipe yields Llama-Fin that outperforms all considered baselines, including large open models at the 70B scale and proprietary models like GPT-4o, on tasks that are similar (yet unseen) to the training data. Even on novel tasks that were never encountered in training, Llama-Fin remains competitive and consistently outperforms its base model across all identified capabilities. In summary, our key contributions are:
  • Comprehensive guidance for finance-specific post-training, including identification of capabilities, evaluation, data and model recipe design.
  • Systematic exploration on each stage of post-training, with an emphasis on the goals, challenges and effective approaches.
  • Novel preference alignment approach that constructs preference data using on-policy trajectories guided by outcome and process signals.
  • New State-of-the-art financial LLM (Llama-Fin) at the 8b parameter scale based on the above.
@@ 2 Related Work
Finance
LLM
Capabilities Recipe Evaluation
@@
AdaptLLM Concept CPT CPT: Financial text + heuristic QAs constructed from the text Financial tasks + Direct answer
PIXIU Task IT IT: Financial tasks Financial tasks + Direct answer
FinLLM Concept, Task CPT → IT CPT: Financial text + Fineweb; IT: Filtered Financial tasks Financial tasks + Direct answer
FinTral Concept, Task CPT → IT → PA CPT: Financial text; IT: Financial tasks; PA: Outcome signal only Financial tasks + Direct answer
@@
Finance LLMs Table 1 summarizes popular finance-specific LLMs developed through domain-adaptive post-training. AdaptLLM (Cheng et al., 2024) focuses on CPT and constructs heuristic QA tasks from raw text, but it considers only five financial end tasks. PIXIU (Xie et al., 2023) focus on instruction-following by creating a financial instruction-tuning dataset from diverse open financial tasks and designing a benchmark with nine end tasks for evaluation. FinLLM (Xie et al., 2024a) extends post-training across multiple stages, first performing CPT, then IT, and incorporating multi-modal capabilities via IT. It includes some general-domain data (e.g., FineWeb (Penedo et al., 2024)) but does not explore its impact systematically. Following this line, FinTral (Bhatia et al., 2024) is the only open FinLLM to include PA, where preference labels were given by GPT-4 on the final outcome, considering only coarse-grained signals. It also introduces multi-modality via IT and integrates tool use and retrieval in PA training. Additionally, Palmyra-Fin (Writer, 2024), a recent state-of-the-art FinLLM, reports high performance on finance tasks, particularly CFA exams [^2] https://www.cfainstitute.org/programs/cfa-program/exam, but its training recipe remains undisclosed.
Comparing to FinDaP, none of these models explicitly identify desirable capabilities as we do with FinCap, nor do they systematically explore trade-offs between CPT, IT and PA to develop a more effective training recipe. They also do not incorporate fine-grained process signals in PA to improve reasoning, as we do in FinRec. Additionally, their evaluations lack the broader range of tasks, methods, and similarities, including reasoning tasks and CoT evaluations, that we adopt in FinEval. Finally, unlike Palmyra-Fin, Llama-Fin is fully open-source, ensuring complete transparency in its training recipe, datasets, and evaluation methods, while achieving SoTA in its size category.
@@ 3.1 Core Capabilities (FinCap)
We began by conducting a comprehensive survey of existing work and consulting two financial domain experts: a banking industry advisor and a financial industry product manager. From this, we identified four key fundamental capabilities essential for a finance LLM: understanding domain-specific concepts to process financial language accurately, performing domain-specific tasks to solve real-world problems, reasoning effectively to analyze complex financial data, and following instructions to interact naturally in practical applications. These capabilities are deeply interconnected: reasoning depends on conceptual knowledge, while instruction-following ensures effective communication.
@@ 3.2.1 Model Recipe
Previous studies often de facto treat domain-adaptive post-training as a sequential process involving, or partially involving, CPT, IT, and PA. However, our experiments with LLaMA3-8B-Inst show key trade-offs among these stages (App. B). While CPT is effective at introducing domain concepts, it often leads to significant forgetting of general concepts and instruction-following capabilities. In contrast, IT strengthens instruction-following capabilities and introduces domain-specific tasks with minimal forgetting. IT alone however struggles with task generalization. PA is effective for learning reasoning but depends heavily on high-quality preference data, which can be difficult to synthesize. To address these limitations, we propose a joint CPT+IT approach, resulting in CPT+IT checkpoint. Subsequently, PA is performed with a novel trajectory collection method that provides fine-grained supervision signals.
@@
Joint continual pre-training and instruction-tuning (CPT + IT). In this stage, the goal is to learn domain-specific knowledge while maintaining general capabilities, such as instruction-following. It is well known that CPT can adapt the LLM to learn domain-specific concepts while IT can help learn the domain-specific and instruction-following tasks Ke et al. (2022); Wei et al. (2022). Typically CPT involves next-token prediction without masking any context tokens, and IT involves next-token prediction with instructions masked out; thus training them sequentially from an instruction-tuned LLM naturally leads to forgetting general capabilities, including instruction-following. Intuitively, if the loss function incorporates both CPT and IT, forgetting can be largely mitigated Scialom et al. (2022). [^3] This is akin to ‘replay’ method in continual learning. To achieve this, we mix CPT and IT data, effectively performing joint optimization, as the only difference between the two is whether the instruction is masked. This approach also facilitates knowledge transfer, as CPT helps the model learn domain knowledge, which can be leveraged by IT training. More importantly, since concepts learned from CPT are often inherently more generalizable due to the shared nature of concepts across tasks, jointly training CPT and IT can improve generalization without require exposure to a diverse range of tasks , which is often impractical in certain domains, particularly long-tail ones. Since CPT datasets are typically much larger than IT datasets, we downsample CPT data to match the size of IT data, allowing for effective joint training.
@@
Improving reasoning with preference alignment. CPT+IT improves capabilities such as in general and domain-specific concepts, tasks and IF/Chat. However, we find that the resulting model lacks in its reasoning capability, especially when it comes to complex reasoning like solving problems in CFA exams, where it is important to make each reasoning step correct. We use PA for this, which trains the model to assign higher probability mass to better generations, and has been shown to be effective in enhancing LLM reasoning capabilities (Lambert et al., 2024; Jiao et al., 2024). Specifically, we employ Direct Preference Optimization or DPO (Rafailov et al., 2023), which allows the model to learn from both positive and negative examples, providing a richer learning signal compared to SFT. We synthetically generate such data from the on-policy model, i.e., the jointly trained CPT+IT checkpoint, as it has shown the strongest performance in preliminary experiments (Appx. B.3). We propose a novel trajectory collection method that provides fine-grained step-level supervision signals (§3.2.2)
@@ 3.2.2 Data Recipe
Mixture of in-domain and general-domain data. Most existing finance LLMs rely exclusively on in-domain data in post-training with the exception of FinLLM, which uses general domain data in CPT (see Table 1). Intuitively, this exclusive reliance on in-domain data can lead to forgetting of general knowledge in the original pre-trained LLM. To understand how the forgetting happens across different stages, we conduct ablations by constructing three versions of data for each training stage: in-domain, general-domain, and a mixture of both. Experiments (App. B) show that the impact of forgetting decreases progressively from CPT to IT to PA, with CPT experiencing the most severe forgetting and PA the least. Guided by these findings, we adopt a mix of in- and general-domain data for CPT+IT training to maximize both specialization and retention of essential general knowledge.
@@
Figure 2: An overview of the proposed final answer preference (FAP) and stepwise corrective preference (SCP). In FAP, we collect trajectories from the GenRM by evaluating the entire solution. In SCP, we collect trajectories from the GenRM, by identifying and correcting the first erroneous step.
@@
Preference data construction for reasoning. Existing training methods to improve reasoning primarily rely on outcome-based rewards, which provide sparse supervision and do not guide intermediate reasoning steps. At the same time, stepwise reward models can be computationally expensive if applied at every step. To strike a balance, we employ a Generative Reward Model (GenRM) and design Final Answer Preference (FAP) to efficiently collect preference signals at the final answer level (outcome reward), while also collecting Stepwise Corrective Preference (SCP) at the reasoning step level (process reward) by asking the GenRM to identify and correct the first erroneous step. By combining these two complementary strategies, our PA provides stronger supervision signals for reasoning improvements while maintaining efficiency, making it particularly suitable for domains like finance, where both accuracy and efficiency are critical. Figure 2 illustrates the proposed method.
@@
$\bullet$ Final Answer Preference (FAP). Given a prompt and a model generated solution, we prompt the GenRM to give a holistic judgment for the entire solution using a single “Yes” or “No” token. We then use the correct solutions as chosen samples and the incorrect solutions as rejected samples.
@@
$\bullet$ Stepwise Corrective Preference (SCP). Since reasoning could be complex (e.g., CFA exams) and process rewards have been shown to be more effective in such cases (Lightman et al., 2024), we further leverage the GenRM to provide step-level signals. Instead of requesting rewards at each step, which has been shown to be unnecessary in Lightman et al. (2024); Luo et al. (2024), we prompt the GenRM to identify the first erroneous step and ask it to provide a correction for that step. Using this correction, we construct a preference data sample. The input prompt is formed by concatenating the original question, the candidate reasoning steps up to the first error, and a follow-up question framed as “What is the next step?”. The chosen response of this preference sample is the newly-obtained corrected step, while the original first erroneous step is deemed as rejected response. This approach produces trajectories that focus on predicting the correct next step given a reasoning prefix, unlike FAP, which requires a prediction of the entire reasoning trajectory (see App. H for prompt details).
@@ 3.3 FinDaP Training Data (FinTrain)
In FinTrain, we carefully balance the trade-off between quality and quantity of the training data at each stage. Specifically, in CPT, we leverage available general-domain supervised data, like NaturalInstrution Mishra et al. (2022). Since such data has been carefully curated and cleaned for labeling, they maintain good quality. For quantity and diversity, which is essential for learning new domain knowledge during CPT, we collect large-scale, diverse data from relevant sources, including 70 financial websites and books covering 12 financial topics, like CFA exam preparation materials. We further use a strong LLM to filter out low-quality tasks based on an additive scale prompt Yuan et al. (2024). This results in approximately 6B tokens.
For IT, to promote diversity, we conduct a broad survey and source general, financial, instruction-following, and reasoning datasets from public datasets. We also include large open QA datasets like FinQA Chen et al. (2021). To ensure quality, we prioritize datasets that shown to perform well in the literature, like UltraChat Ding et al. (2023). We also incorporate exercises or demonstrations from books that often contain human-written CoT. The final IT dataset consists of $\sim$3M prompts.
For PA, we use CFA preparation materials as a representative source for in-domain reasoning as they cover diverse financial scenarios, emphasize complex reasoning, and, most importantly, are derived from real-world exams. We construct preference data with FAP and SCP introduced in §3.2.2. The final PA dataset consists of about 32K prompts. Additional details are given in App. D.
@@ 3.4 FinDaP Evaluation (FinEval)
Our evaluation framework FinEval is designed to systematically assess model performance across unseen tasks. Unlike prior studies that rely on a narrow set of domain-specific tasks, FinEval categorizes tasks by similarity (similar vs. novel tasks), domain specificity (general vs. domain-specific vs. reasoning tasks), and evaluation methods (direct-answer vs. chain-of-thought). By structuring evaluations along these dimensions, FinEval consists of 35 tasks and can serve as a comprehensive benchmark for the expected capabilities going beyond simple task-based evaluation. We took extra care to ensure that FinEval does not duplicate any samples from FinTrain: the 10-gram contamination rate is only 0.003%, indicating minimal overlap (see App. A). We provide details about the evaluation tasks and methods in App. E.